Solve equation for x.
1/27=3^2x

Answer by nerdybill(5412)

1/27=3^2x
Take log3 (log base 3) of both sides:
log3(1/27) = x
log3(1) - log(327) = x
log3(3^0) - log3(3^3) = x
0-3 = x
-3 = x